### v1

# Communication Optimization for Distributed Training: Architecture, Advances, and Opportunities

Abstract
The past few years have witnessed the flourishing of large-scale deep neural network models with ever-growing parameter numbers. Training such large-scale models typically requires massive memory and computing resources that exceed those of a single GPU, necessitating distributed training. As GPU performance has rapidly evolved in recent years, computation time has shrunk, thereby increasing the proportion of communication in the overall training time. Therefore, optimizing communication for distributed training has become an urgent issue. In this article, we briefly introduce the general architecture of distributed deep neural network training and analyze relationships among Parallelization Strategy, Collective Communication Library, and Network from the perspective of communication optimization, which forms a three-layer paradigm. We then review current representative research advances with this three-layer paradigm. We find that layers in the current three-layer paradigm are relatively independent, but there is a rich design space for cross-layer collaborative optimization in distributed training scenarios. Therefore, we further advocate a communication-efficient five-layer paradigm underlining opportunities for collaboration designs and look forward to the perspectives of “Vertical”, “Horizontal”, “Intra-Inter” and “Host-Net” collaboration designs. We hope this article can shed some light on future research on communication optimization for distributed training.
Index Terms:
Deep Neural Network, Distributed Training, Parallelization Strategy, Collective Communication Library, Network Protocols and Topologies.I Introduction
Large-scale deep neural network (DNN) models (“large models” for short) have become ubiquitous, and their capabilities have significantly advanced in recent years. The latest prominent large models such as GPT111https://openai.com/gpt-4, LLaMA222https://llama.meta.com/llama2, and GLM333https://open.bigmodel.cn demonstrated unprecedented performance, enabling significant changes in production and lifestyle.
The rapid development of large models inevitably led to a significant increase in the scale of parameters and training data. Due to the limited memory and computing power of a single GPU (for simplicity, we use GPU to signify other AI hardware such as TPU and NPU), training large models with a single GPU has already been a thing of the past. For instance, even merely considering the computation time, training a GPT-3 model with 175 billion parameters on a single Nvidia V100 GPU would require approximately 288 years [1]. Consequently, training large models with multiple GPUs is natural and inevitable, creating a new demand for large-scale, high-performance GPU clusters.
However, enhancing GPU performance and enlarging cluster size do not necessarily lead to linear performance improvement of distributed deep neural network training (“distributed training” for short) systems, as one might naturally expect. This is because the time overhead of distributed training comes not only from computation but also from communication. When the time spent on computation is reduced, the communication time is exposed more and gradually becomes a bottleneck. In fact, the communication tasks account for up to 60% of a DNN training iteration time in Meta’s production environment [2]. Optimizing communication can significantly reduce the overall training time, so there is an urgent need to enhance communication efficiency for distributed training.
Systematically optimizing distributed training is not a straightforward task. Distributed deep neural network training architecture consists of hardware and software parts. The software part includes parallelization strategy, deep learning library, collective communication library (CCL), and network protocols, while the hardware part comprises CPUs, GPUs, RAM, I/O, and network infrastructure. Attaining efficiency in distributed training systems necessitates effective collaboration among various system components[3, 4].
Many components of distributed training architecture are closely related to communication. For instance, parallelization strategy determines communication demand, CCL generates communication traffic, and network affects the efficiency of performing communication tasks. To sum up, there is a “Parallelization Strategy, CCL, and Network” three-layer paradigm of communication optimization for distributed training.
There are many research advances optimizing communication within the current three-layer paradigm. For example, PTD-P [1] uses a novel interleaved pipeline parallelism optimization scheme to overlap communication and computation as much as possible. TACCL [5] generates communication primitive algorithms tailored to specific training tasks and topology to enhance efficiency. TopoOpt [2] leverages the reconfigurability of optical switches to optimize topologies and parallel strategies collaboratively, providing customized topology for efficient communication.
Different from general high-performance computing (HPC) scenarios, distributed training has its own characteristics. For example, communication traffic of distributed training exhibits periodic repetition, and the traffic pattern of each training iteration is relatively consistent [6]. Moreover, most communication tasks in distributed training are pre-determined, which contrasts significantly with traditional tasks where traffic flows arrive stochastically. Meanwhile, the current three-layer paradigm has its shortcomings. Each layer of the paradigm is relatively independent, making it difficult for them to cooperate for communication optimization. Therefore, we advocate a communication-efficient five-layer paradigm, which is detailed in Section IV, aiming to achieve cross-layer collaboration design with (logical) schedulers as middleware. We prospect optimization opportunities from four perspectives: “Vertical”, “Horizontal”, “Intra-Inter”, and “Host-Net” co-design.
The rest of this article is arranged as follows: Section II presents the general architecture of distributed training and the current three-layer paradigm from the perspective of communication optimization. Section III reviews representative research advances on communication optimization for distributed training. Section IV advocates a communication-efficient five-layer paradigm, looking into the cross-layer collaborative design opportunities from four promising research directions. Finally is the conclusion.
II Architecture
This section first provides an overview of the distributed deep neural network training architecture, followed by an in-depth introduction of three components pertinent to communication optimization: parallelization strategy, collective communication library, and network. Finally, the relationships between these three components and communication optimization are elaborated.
II-A Overview
The general architecture of the distributed deep neural network training system is shown in Fig. 3. Top of the architecture is a large deep neural network model to train. As analyzed in Section I, a deep learning model needs distributed training when it is too large to fit into the memory or too slow to train with a single GPU.
Parallelization strategy is an essential part of distributed training deployment, determining how a model is partitioned for distributed training. Commonly used parallelization strategies are detailed in Section II-B.
Deep learning models are often implemented with the deep learning library (such as TensorFlow444https://www.tensorflow.org and PyTorch555https://pytorch.org) , which generates execution task graphs including computing tasks and corresponding collective communication tasks. Communication compression techniques such as gradient quantization are often more relevant to training methods and are not detailed in this article due to the limited space.
The deep learning library often invokes Collective Communication Library to implement collective communication tasks to transmit activations (in forward propagation) or synchronize gradients (in backward propagation) between different GPUs. Commonly used collective communication primitives implemented in CCL are detailed in Section II-C.
The collective communication primitives generate actual communication data traffic, which is injected into the underlying network. A variety of network protocols and topologies for distributed training are detailed in Section II-D. As the end-host of the underlying network, a computing node is a supercomputer equipped with high-performance GPUs and CPUs, a large RAM, and high-speed I/O. These hardware components also need to collaborate efficiently.
Parallelization strategy, CCL, and Network are three critical components in the architecture that affect communication efficiency. The interplay among them and their impacts on communication efficiency are analyzed in Section II-E.
II-B Common Parallelization Strategies
Commonly used parallelization strategies include data parallelism, model parallelism (primarily including pipeline parallelism and tensor parallelism), and emerging Mixture-of-Expert (MoE) parallelism, as illustrated by Fig. 3, where each rectangle refers to a GPU node and each round rectangle refers to (part of) a model.
Data parallelism is one of the most commonly used parallelization strategies whose basic idea is to distribute multiple copies of a model to different GPUs. Each GPU receives a subset of training data (mini-batch) during training, and gradients are averaged to update global parameters at the end of each iteration. Model parallelism, including pipeline and tensor parallelism, means splitting the model onto different GPUs. Pipeline parallelism allocates different layers of the model to different GPUs, so there is mainly point-to-point communication between layers (GPUs). Tensor parallelism splits the same layer of the model onto different GPUs and uses distributed matrix computing techniques for collaboration [7], which is a communication-intensive operation. MoE parallelism involves dividing a portion of a model into multiple expert components. Each expert specializes in a specific task domain and is allocated to a certain GPU.
In practice, the above parallelization strategies are often not used independently, but in a hybrid manner. For example, we can first divide the model by layers and split each layer, then distribute the divided model to multiple sets of GPUs, thereby achieving pipeline-tensor-data three-dimensional hybrid parallelism [1]. MoE parallelism is another example of hybrid parallelism. The idea of data parallelism is implicit in MoE parallelism, and the expert model itself can also be split into multiple GPUs for parallel computing.
II-C Common Collective Communication Primitives
Commonly used collective communication primitives implemented in CCL for distributed training are illustrated in Fig. 3, where each rectangle represents a GPU node and each rounded rectangle represents a data chunk.
Broadcast distribute data from a particular node to all other nodes, which can be used in data parallelism and some model parallelisms [8]. All-Gather is a many-to-many collective communication primitive where data from different nodes are distributed to all nodes. All-to-All transmit data among various nodes, such as data distribution in MoE parallelism [9, 10]. All-Reduce is a sum operation of the corresponding data chunk of each node, as shown in Fig. 3(d), where each grey rounded rectangle represents the aggregated results of the data chunks at the corresponding position of each GPU node. Common All-Reduce scenarios include data parallelism and some model parallelism such as Megatron-lm [7].
II-D Underlying Network
Protocols and topologies are two main factors that affect network performance. Besides general TCP/IP protocol, distributed training often uses RDMA (such as RoCE or Infiniband) for less overhead and higher bandwidth.
The traffic flow within the network is also closely related to the network topology. Common network topologies used for distributed training include Fat-tree and its variants, Torus, as well as Ring and Full-Mesh topologies. These topologies can be combined according to practical needs. For example, the NVLink topology of Nvidia’s DGX-1666https://images.nvidia.com/content/pdf/dgx1-system-architecture-whitepaper1.pdf is a combination of Ring and Full-mesh.
II-E Communication Paradigm in the Architecture
Parallelization strategy, CCL, and network form a three-layer paradigm of communication optimization for distributed training, as shown by the colored rectangles in Fig. 3. All of them directly or indirectly affect the communication performance of the training process. Different parallelization strategies determine the primarily used collective communication primitives in task graphs, which affects the traffic pattern. Various implementations of CCL directly affect the actual network traffic. The same traffic demand often exhibits different performance under different network infrastructures. The topologies designed for distributed training are closely related to the algorithm implementation of collective communication primitives. For example, the execution process of the Ring-based All-Reduce algorithm results in a ring communication mode. That is, each node transmits data to its logical neighbors while all the nodes in the communicator (a communicator is a set of nodes used to implement a collective communication task) form a ring. Nvidia’s DGX-16 and Google’s Torus [4] topology both contain numerous ring structures, which are well-suited to satisfy the communication needs of Ring-based All-Reduce.
However, layers are relatively independent in the current three-layer communication paradigm. Although some methods introduced in Section III involve cross-layer design, collaborative design has not yet become mainstream in general, which limits the end-to-end training performance [6]. Therefore, we advocate a communication-efficient five-layer paradigm with additional (logical) schedulers (dashed boxes in Fig. 3) over the current architecture. The inter-layer schedulers imply collaborative design, which is elaborated in Section IV.
III Overview of Recent Advances
Communication optimization in distributed training encompasses numerous aspects. We delve into three of the most pertinent elements: parallelization strategy, CCL, and network, which make up the current three-layer communication paradigm. We use exposed communication to denote the communication time that cannot overlap with computation time. In this section, our focus is confined to the contributions of a specific study pertaining to communication optimization, that is, how to reduce exposed communication in distributed training. Table I provides an overview of the research advances with some primary concerns.
Layer |
Main focus |
How to reduce exposed communication |
Cross-layer co-design |
Time for decision |
Implementation |
Method |
Beneficial effect |
Megatron-lm [7] |
Model Parallelism |
Efficient intra-layer model parallel strategy |
An efficient tensor partitioning strategy |
Reducing the amount of traffic by removing a synchronization point |
Offline |
Insert a few communication operations in naive pytorch |
AlpaComm [8] |
Addressing “cross-mesh resharding” problem in model parallelism |
Broadcast based resharding method and eager 1F1B pipeline schedule |
Reducing inter-node traffic and overlap communication with computation |
With CCL |
On/Offline |
900 lines of C++ and 2.5K lines of python |
PTD-P [1] |
Hybrid Parallelism |
Combination of pipeline, tensor and data parallelisms |
A interleaved pipelining schedule |
Overlapping communication and computation in pipeline parallelism |
Offline |
An extension to Megatron-LM codebase |
Lina [9] |
MoE Parallelism |
All-to-All bottleneck in MoE parallelism |
Priority scheduling All-to-All with All-Reduce micro-ops |
Overlapping All-Reduce communication with computation |
With CCL |
Offline |
Based on DeepSpeed MoE with 7500 lines of code |
Janus [10] |
A data-centric paradigm for MoE parallelism |
Move experts to reduce All-to-All traffic when benificial and a provident prefetch strategy |
Reducing All-to-All traffic and overlap communication with computation |
On/Offline |
A plugin of 4.3K lines of code in PyTorch |
NCCL |
Collective Communi- cation Library (CCL) |
Multi-GPU and multi-node communication primitives optimized for NVIDIA hardware |
Select communication primitive algorithms such as DBTree and Ring for All-Reduce, etc. |
Mainly accelerating All-Reduce communication |
For particular hardware |
On/Offline |
Integrated in Deep learning frameworks such as PyTorch and TensorFlow. |
Blink [11] |
Synthesize collective communication algorithms for underlying topology |
Packing spanning trees with MWU and ILP algorithms |
Overlap flows within primitives to promote link utilization |
For topology |
Online |
Seamlessly plugged into pytorch and tensorflow |
SCCL [12] |
Synthesize collective communication algorithms for underlying topology |
Synthesizing pareto-optimal algorithms with SMT encoding |
Overlap flows within primitives to promote link utilization |
For topology |
Offline |
Evaluated on different 8-GPU systems such as Nvidia DXG-1 and Gigabyte Z52 |
TACCL [5] |
Synthesize collective communication algorithms for topology and training task |
Routing (MILP), ordering (Greedy algorithm) and scheduling (MILP) with communication sketches |
Overlap flows within primitives to promote link utilization |
For topology |
Offline |
Extension of NCCL with TACCL runtime as interpreter |
SYNDI-CATE [13] |
Co-optimize collective communication algorithms with runtime schedule |
MCMC Search with termination |
Overlap primitives to promote link utilization |
With execution plan |
On/Offline |
UCC library interface (pytorch integrated) |
TPUv4 [4] |
Network |
Google’s third supercomputer specific for ML |
Architectural |
Up to 4096 chips connected by OCSs |
TopoOpt [2] |
Co-optimizing computation, communication and network topology |
A group theory-inspired algorithm called TotientPerms |
The best network topology and routing plan, together with a parallelization strategy |
With parallelization strategy |
Offline |
A 12-node prototype with FlexFlow’s training engine and modification to NCCL |
III-A Parallelization Strategy
The communication overhead in distributed training mainly arises from parallelization. Different implementations of parallelization strategies affect the communication pattern and network traffic during the model training process [2].
The most straightforward and most widely used parallelization strategy is data parallelism. As mentioned in Section II, multiple copies of the same model are replicated across different nodes in data parallelism. Each iteration requires all nodes to synchronize gradients, which are used for global parameter updates. Consequently, data parallelism results in a significant amount of All-Reduce communication.
To train a large-scale model that can’t fit into the memory of a single GPU, researchers introduced model parallelism, which includes pipeline and tensor parallelism. Tensor parallelism divides models within a layer and requires intensive communication in training. Researchers from Nvidia proposed an efficient “intra-layer parallelism” approach called Megatron-lm [7] to reduce the communication traffic in tensor parallelism. Pipeline parallelism necessitates only the transfer of activations/gradients between adjacent nodes in the logical pipeline, which is a point-to-point communication pattern that significantly reduces the communication cost compared to data parallelism and tensor parallelism. Because overlapping communication and computation is hard to achieve in naive pipeline parallelism, an interleaved pipeline scheduling strategy [1] is proposed where each device is assigned multiple data chunks to gain the chance of communication-computation overlap. However, pipeline parallelism, as an “inter-layer parallelism” approach, is not without its flaws. It faces issues such as “synchronous/asynchronous” parameter updates and pipeline bubble problems, which can adversely affect the accuracy of the model and the efficiency of training [7].
Various parallelization strategies are often used in conjunction. For example, PTD-P [1] not only combines Megatron-lm’s Tensor parallelism and pipeline parallelism but also introduces data parallelism to scale to thousands of GPUs. AlpaComm [8] identifies the “cross-mesh resharding” problem that arises when combining tensor and pipeline parallelism and proposes a communication optimization method based on broadcasting and an “overlapping-friendly” pipeline scheduling scheme to accelerate end-to-end training.
As an emerging model structure, the MoE model is gradually gaining popularity due to its fast training/inference speed. The training process of the MoE model selects several appropriate experts for a data sample, which inherently has good compatibility with parallelization. MoE parallelism can be achieved by simply dispersing the “experts” on different GPUs, introducing All-to-All traffic when data is distributed between different GPUs. Additionally, the model parameters of the non-expert part are globally shared in MoE parallelism, so there will also be All-Reduce traffic. Lina [9] uses a scheduling strategy that prioritizes All-to-All traffic and further splits the All-Reduce tasks to overlap with computation as much as possible, thus reducing exposed communication and achieving training acceleration. Janus [10] studies MoE parallelism acceleration from a new perspective. Since the size of expert models and training datasets varies, Janus innovatively proposes a “data-centric” model that “moves experts instead of data”. This significantly reduces the amount of communication during the training process under certain conditions (e.g., when the scale of experts is smaller than the scale of data), thus accelerating the training process.
III-B Collective Communication Library
Communication tasks of distributed training are often implemented by CCL. The current predominant NVIDIA Collective Communication Library (NCCL)777https://developer.nvidia.com/nccl dynamically selects established primitive algorithms based on different situations. For instance, NCCL selects the Ring and Double Binary Tree algorithms for All-Reduce 888https://developer.nvidia.com/blog/massively-scale-deep-learning-training-nccl-2-4 for different workloads and topologies to accelerate the training process. However, NCCL’s optimization for specific hardware underscores the need for a more universal and adaptable CCL to speed up distributed training across different hardware and topologies. Therefore, many researchers took different approaches to implement communication primitives more efficiently.
For example, Blink [11] dynamically generates optimal communication primitives by packing spanning trees rather than writing and optimizing them manually. Blink uses integer linear programming (ILP) to reduce the number of trees and improve performance, and also leverages heterogeneous communication channels to utilize full bandwidth. At the time, Blink achieved up to 40% speedup compared to NCCL in end-to-end DNN training iterations.
SCCL [12] takes a more unified approach to achieve efficiency: automatically synthesizing high-performance communication primitives for a given topology. To achieve this, SCCL designs a cost model to evaluate the latency and bandwidth cost of algorithms and then searches for algorithms on the Pareto frontier. However, SCCL also faces problems of high complexity. The communication primitive algorithm generation problem is encoded into a Mixed Integer Linear Program (MILP), which is unfortunately NP-hard.
To reduce the complexity of generating collective communication algorithms, TACCL [5] is a representative work that introduces a “human-in-the-loop” approach, whose paradigm is shown in Fig. 4. TACCL incorporates high-level inputs from an algorithm designer, such as logical topologies, switch hyper-edges, and algorithm symmetry, to efficiently synthesize collective communication algorithms for heterogeneous topologies. These human inputs greatly constrain the search space of algorithms, reducing search time to an acceptable amount, in most cases from seconds to a few minutes. Compared with NCCL, TACCL achieves up to 2.36x end-to-end training speedup for BERT and 1.94x for Transformer-XL.
Different from previous methods focused on speeding up single primitives, SYNDICATE [13] proposes a novel abstraction to overlap primitives in a fine-grained pattern, breaking large communication operations into smaller units, thus allowing higher flexibility. SYNDICATE then develops a joint optimizer based on Markov Chain Monte Carlo (MCMC) search to optimize the joint action of the control plane and the data plane.
III-C Network
Distributed training focuses on GPU-to-GPU communication, which needs to consider intra/inter-host communication protocols. Common intra-host communication protocols include the general PCIe protocol and Nvidia’s NVLink/NVSwitch. Inter-host communication protocols include traditional TCP/IP and more efficient RDMA (e.g., RoCE and Infiniband).
Based on certain communication protocols, the underlying network topology also significantly influences the communication efficiency of distributed training. For instance, as mentioned in Section II, Nvidia utilizes a Ring and fully connected Mesh topology in their DXG series, and Google employs a 3D-Torus topology in their TPUv4 [4]. These topologies contain a large number of ring structures, which enhance the efficiency of Ring-based All-Reduce and Point-to-Point communications (such as pipeline parallelism or broadcast-based hybrid parallelism [8]). However, static topologies can rarely adapt to dynamically changing job requirements, leading researchers to propose TopoOpt [2]. TopoOpt takes advantage of the reconfigurability of optical switches, enabling a scheme that co-optimizes the topology and parallelization strategy. Nevertheless, current optical switching technology still has issues of high reconfiguration latency, making it challenging to perform dynamic topology adjustments between iterations. As such, TopoOpt remains a relatively coarse-grained optimization, that is, reconfiguring the interconnection between training servers of each job only before the job starts and keeps the topology until the training is complete.
Anyway, research on communication optimization for distributed training from a network perspective is still insufficient, which is prospected in Section IV.
III-D Collaboration Design in Current Advances
As observed in Table I, the research advances mentioned above are primarily limited to optimizations within a layer, with less cross-layer co-design. Current advances involving cross-layer optimization can be categorized into two types. The first type is passive collaborative design. For instance, AlpaComm [8] and Lina [9] optimize communication for a specific parallelization strategy, necessitating modifications to communication primitives. Another example is that generative CCL (such as Blink [11], SCCL [12] and TACCL [5]) has to be aware of network topologies because their main significances are to customize collective communication algorithms for specific topologies. The second type of work involves proactive attempts at collaborative optimization yet remains exploratory. For example, SYNDICATE [13] jointly optimizes the implementation and scheduling of communication primitives, allowing parallel execution of primitives. TopoOpt [2] collaboratively optimizes network topology and parallelization strategies, obtaining optimized combination strategies through alternating optimization techniques.
In summary, collaborative communication optimization in distributed training is still a domain that needs further exploration. Section IV details the importance of co-design and potential research opportunities.
IV Opportunities
As noted in Section III, collaborative design has not received widespread attention in current research advances. However, different from general tasks in data centers, the communication traffic patterns in distributed training are known in most cases [6, 14], providing motivations and opportunities for cross-layer collaboration. Furthermore, the optimization objectives for communication tasks are no longer the general-purpose flow complete time (FCT) but rather the job complete time (JCT). In this context, communication tasks are merely a component of the entire distributed training process. Optimizing JCT necessitates considering the dependency between communication and computation instead of just optimizing each communication task. Communication tasks should be optimized to minimize computation blocking time, which can only be achieved through “cross-layer” scheduling to meet this new optimization objective.
To this end, we advocate a communication-efficient five-layer paradigm with (logical) schedulers as the middleware, as illustrated in Fig. 5. The schedulers not only allow lower-layer facilities to perceive the requirements of higher-layer applications but also enable higher-layer applications to be aware of the capabilities of lower-layer facilities. We refer to this as “Vertical” co-design. The schedulers can also perform global scheduling across multiple jobs, achieving “Horizontal” collaborative optimization. Considering the heterogeneous topologies of distributed training and the development of emerging technologies such as in-network aggregation, we also provide a prospect for “Intra-Inter” and “Host-Net” co-design.
“Vertical” co-design: Refers to exchanging necessary information between layers to optimize communication efficiency. For example, CCL can acquire global traffic demand from the parallelization strategy and generate the optimal communication primitive algorithms in conjunction with the network state. Furthermore, cross-layer schedulers can be constructed to strategically schedule communication tasks based on traffic load or optimize task graphs according to the traffic patterns of relevant communication primitives. Echelon [14] is a flow scheduling work that embodies “Vertical” co-design, where novel communication optimization objectives are designed based on the dependency between communication and computation under various parallelization strategies, aiming to optimize communication towards minimizing JCT.
Another example of “Vertical” co-design is that network can also use traffic demand for optimization in distributed training. As the general infrastructure, the network is often unaware of traffic demand. This is natural in common data center scenarios, where the network can only implement passive adjustments such as RTT-based congestion control due to the random nature of flow arrivals. However, traffic patterns are often predictable in distributed training scenarios[6]. Consequently, the network itself can also proactively adjust by taking traffic demands of high-layer training tasks into account, thus enhancing goodput. This in turn reduces communication time and accelerates the distributed training process.
“Horizontal” co-design: Primarily focuses on the coordination among different training jobs or various communication tasks. Within a single training job, the traffic from multiple concurrent collective communication primitives competes for network resources. Similarly, in the multi-job scenario, the traffic from numerous concurrent training jobs also competes for network resources. Consequently, optimization strategies tailored to a single communication primitive or training job may suffer varying degrees of inefficacy. Therefore, it is imperative to employ horizontal co-design at different granularities to fully utilize network resources, ultimately reducing JCT of training tasks. CASSINI [6] attempts horizontal co-design through the perspective of multi-task scheduling. It analyses the consequence of non-coordinated job scheduling in multi-job scenarios, highlighting the potential traffic conflicts arising from the periodic patterns of network resource utilization by different training jobs. Subsequently, CASSINI introduces a design for “staggering peak” scheduling for different jobs, achieving high utilization of network resources, thus enhancing the performance of multiple training jobs.
“Intra-Inter” co-design: Refers to the coordinated utilization of multidimensional network resources, primarily including the “intra-host” and “inter-host” network. The current networks used in distributed training exhibit substantial heterogeneity. The interconnection patterns between different GPUs within the same host include PCIe or NVLink, while those between different hosts include RoCE or Infiniband, etc. Given that GPU (rather than host) is the minimum unit in the communication tasks of distributed training, two modes of GPU-to-GPU communication exist: within the same host and across different hosts. In this context, the issue of heterogeneous link bandwidths within and between hosts is inevitably confronted. An example of “Intra-Inter” co-design is that TACCL [5] utilizes a profiling approach to perceive the characteristics of heterogeneous links, which serves as a basis for optimizing communication primitive algorithms.
“Host-Net” co-design: Primarily refers to the collaboration of computing resources between the end-host and programmable switches within a network. In recent years, with the development of programmable switch technology, in-network aggregation has gradually become a new computing paradigm. In-network aggregation not only fully utilizes the computing capabilities of programmable switches within the network but also reduces the traffic running in the network, achieving multiple purposes in one stroke. However, since existing protocols are designed for end-to-end communication, implementing in-network aggregation usually requires the support of new protocols. For instance, ATP [15] is an in-network aggregation transmission protocol designed for multi-tenant scenarios, which can make full use of the computing capabilities of multi-level programmable switches and degrade to host aggregation when the in-network computing resources are exhausted. This approach also reduces the network traffic for a collective communication task. ATP embodies the concept of “Host-Net” co-design and represents an exemplary work of in-network computing in the field of distributed training. Nevertheless, in-network aggregation has not been widely applied in the production environment due to its complexity. “Host-Net” co-design remains a direction worthy of long-term research.
V Conclusion
Due to the dramatic increase in the number of parameters of large-scale deep neural network models, it is imperative to develop effective communication optimization methods in distributed training. In this article, we first provide a general distributed training architecture and the current three-layer paradigm of communication optimization consisting of parallelization strategy, CCL, and network. We then review the latest advances in the current three-layer paradigm. Furthermore, we analyze the shortcomings of the current paradigm and advocate a communication-efficient five-layer paradigm, highlighting great opportunities for “cross-layer” co-design. We hope to shed some light on future research on communication optimization for distributed training.

### v2

@@
 The past few years have witnessed the flourishing of large-scale deep neural network models with ever-growing parameter numbers.
-Training such large-scale models typically requires massive memory and computing resources that exceed those of a single GPU, necessitating distributed training.
-As GPU performance has rapidly evolved in recent years, computation time has shrunk, thereby increasing the proportion of communication in the overall training time.
-Therefore, optimizing communication for distributed training has become an urgent issue.
+Training such large-scale models typically requires massive memory and computing resources, necessitating distributed training.
+As GPU performance has rapidly evolved in recent years, computation time has shrunk, making communication a larger portion of the overall training time.
+Consequently, optimizing communication for distributed training has become crucial.
 In this article, we briefly introduce the general architecture of distributed deep neural network training and analyze relationships among Parallelization Strategy, Collective Communication Library, and Network from the perspective of communication optimization, which forms a three-layer paradigm.
-We then review current representative research advances with this three-layer paradigm.
-We find that layers in the current three-layer paradigm are relatively independent, but there is a rich design space for cross-layer collaborative optimization in distributed training scenarios.
-Therefore, we further advocate a communication-efficient five-layer paradigm underlining opportunities for collaboration designs and look forward to the perspectives of “Vertical”, “Horizontal”, “Intra-Inter” and “Host-Net” collaboration designs.
+We then review current representative research advances within this three-layer paradigm.
+We find that layers in the current three-layer paradigm are relatively independent and there is a rich design space for cross-layer collaborative optimization in distributed training scenarios.
+Therefore, we advocate “Vertical” and “Horizontal” co-designs which extend the three-layer paradigm to a five-layer paradigm.
+We also advocate “Intra-Inter” and “Host-Net” co-designs to further utilize the potential of heterogeneous resources.
 We hope this article can shed some light on future research on communication optimization for distributed training.
@@
 Large-scale deep neural network (DNN) models (“large models” for short) have become ubiquitous, and their capabilities have significantly advanced in recent years.
-The latest prominent large models such as GPT111https://openai.com/gpt-4, LLaMA222https://llama.meta.com/llama2, and GLM333https://open.bigmodel.cn demonstrated unprecedented performance, enabling significant changes in production and lifestyle.
-The rapid development of large models inevitably led to a significant increase in the scale of parameters and training data.
-Due to the limited memory and computing power of a single GPU (for simplicity, we use GPU to signify other AI hardware such as TPU and NPU), training large models with a single GPU has already been a thing of the past.
-For instance, even merely considering the computation time, training a GPT-3 model with 175 billion parameters on a single Nvidia V100 GPU would require approximately 288 years [1].
-Consequently, training large models with multiple GPUs is natural and inevitable, creating a new demand for large-scale, high-performance GPU clusters.
-However, enhancing GPU performance and enlarging cluster size do not necessarily lead to linear performance improvement of distributed deep neural network training (“distributed training” for short) systems, as one might naturally expect.
+The latest prominent large models such as GPT111https://openai.com/gpt-4, LLaMA222https://llama.meta.com, and GLM333https://open.bigmodel.cn demonstrated unprecedented performance, enabling significant changes in production and lifestyle.
+The rapid development of large models leads to a significant increase in the scale of parameters and training data.
+Due to the limited memory and computing power of a single GPU, training large models with a single GPU has already been a thing of the past.
+For instance, training a GPT-3 model with 175 billion parameters on a single Nvidia V100 GPU would require approximately 288 years [1].
+Consequently, training large models with multiple GPUs is natural and inevitable.
+This creates a new demand for large-scale, high-performance GPU clusters to accelerate the model training process.
+However, simply enhancing GPU performance and enlarging cluster size do not necessarily lead to linear performance improvement of distributed deep neural network training (“distributed training” for short) systems, as one might naturally expect.
 This is because the time overhead of distributed training comes not only from computation but also from communication.
 When the time spent on computation is reduced, the communication time is exposed more and gradually becomes a bottleneck.
-In fact, the communication tasks account for up to 60% of a DNN training iteration time in Meta’s production environment [2].
-Optimizing communication can significantly reduce the overall training time, so there is an urgent need to enhance communication efficiency for distributed training.
-Systematically optimizing distributed training is not a straightforward task.
-Distributed deep neural network training architecture consists of hardware and software parts.
-The software part includes parallelization strategy, deep learning library, collective communication library (CCL), and network protocols, while the hardware part comprises CPUs, GPUs, RAM, I/O, and network infrastructure.
-Attaining efficiency in distributed training systems necessitates effective collaboration among various system components[3, 4].
-Many components of distributed training architecture are closely related to communication.
-For instance, parallelization strategy determines communication demand, CCL generates communication traffic, and network affects the efficiency of performing communication tasks.
-To sum up, there is a “Parallelization Strategy, CCL, and Network” three-layer paradigm of communication optimization for distributed training.
-There are many research advances optimizing communication within the current three-layer paradigm.
+The communication overhead can account for up to 60% of a DNN training iteration time in Meta’s production environment [2].
+Therefore, there is an urgent need to enhance communication efficiency for distributed training.
+Many components of distributed training systems[3, 4] are closely related to communication.
+For instance, parallelization strategy determines communication demand, collective communication library (CCL) generates communication traffic, and network affects the efficiency of performing communication tasks.
+To sum up, there is a Parallelization Strategy, CCL, and Network three-layer paradigm of communication optimization for distributed training.
+There are many research advances optimizing communication for distributed training in the current three-layer paradigm.
 For example, PTD-P [1] uses a novel interleaved pipeline parallelism optimization scheme to overlap communication and computation as much as possible.
@@
 TopoOpt [2] leverages the reconfigurability of optical switches to optimize topologies and parallel strategies collaboratively, providing customized topology for efficient communication.
-Different from general high-performance computing (HPC) scenarios, distributed training has its own characteristics.
-For example, communication traffic of distributed training exhibits periodic repetition, and the traffic pattern of each training iteration is relatively consistent [6].
-Moreover, most communication tasks in distributed training are pre-determined, which contrasts significantly with traditional tasks where traffic flows arrive stochastically.
-Meanwhile, the current three-layer paradigm has its shortcomings.
-Each layer of the paradigm is relatively independent, making it difficult for them to cooperate for communication optimization.
-Therefore, we advocate a communication-efficient five-layer paradigm, which is detailed in Section IV, aiming to achieve cross-layer collaboration design with (logical) schedulers as middleware.
-We prospect optimization opportunities from four perspectives: “Vertical”, “Horizontal”, “Intra-Inter”, and “Host-Net” co-design.
+Unlike general high-performance computing scenarios, distributed training has distinct characteristics.
+For instance, the communication traffic of each iteration is typically fixed [6] and most of the communication tasks are pre-determined, which contrasts with the stochastic traffic in traditional scenarios.
+However, components in the current communication paradigm often function independently, hindering collaborative optimization for communication.
+In Section IV, we explore research opportunities from a collaborative design perspective to reduce job completion time (JCT) and enhance the efficiency of distributed training systems.
 The rest of this article is arranged as follows: Section II presents the general architecture of distributed training and the current three-layer paradigm from the perspective of communication optimization.
-Section III reviews representative research advances on communication optimization for distributed training.
-Section IV advocates a communication-efficient five-layer paradigm, looking into the cross-layer collaborative design opportunities from four promising research directions.
+Section III reviews representative research advances.
+Section IV illustrates the necessity of collaborative design through a case study and provides an outlook on research opportunities.
 Finally is the conclusion.
 II Architecture
-This section first provides an overview of the distributed deep neural network training architecture, followed by an in-depth introduction of three components pertinent to communication optimization: parallelization strategy, collective communication library, and network.
-Finally, the relationships between these three components and communication optimization are elaborated.
+This section first provides an overview of the distributed training architecture from the perspective of communication, followed by an in-depth introduction of three components pertinent to communication optimization: parallelization strategy, CCL, and network.
+Finally, we summarize the three components as a three-layer paradigm of communication optimization and analyze their impacts on communication efficiency.
 II-A Overview
-The general architecture of the distributed deep neural network training system is shown in Fig.
-3.
-Top of the architecture is a large deep neural network model to train.
-As analyzed in Section I, a deep learning model needs distributed training when it is too large to fit into the memory or too slow to train with a single GPU.
-Parallelization strategy is an essential part of distributed training deployment, determining how a model is partitioned for distributed training.
+From the perspective of communication, the general architecture of the distributed training system is shown in Fig.
+1.
+Top of the architecture is a large deep neural network model that requires distributed training.
+Parallelization strategy is an essential part of distributed training deployment which determines how a model is partitioned for distributed training.
 Commonly used parallelization strategies are detailed in Section II-B.
 Deep learning models are often implemented with the deep learning library (such as TensorFlow444https://www.tensorflow.org and PyTorch555https://pytorch.org) , which generates execution task graphs including computing tasks and corresponding collective communication tasks.
-Communication compression techniques such as gradient quantization are often more relevant to training methods and are not detailed in this article due to the limited space.
-The deep learning library often invokes Collective Communication Library to implement collective communication tasks to transmit activations (in forward propagation) or synchronize gradients (in backward propagation) between different GPUs.
-Commonly used collective communication primitives implemented in CCL are detailed in Section II-C.
-The collective communication primitives generate actual communication data traffic, which is injected into the underlying network.
-A variety of network protocols and topologies for distributed training are detailed in Section II-D.
-As the end-host of the underlying network, a computing node is a supercomputer equipped with high-performance GPUs and CPUs, a large RAM, and high-speed I/O.
-These hardware components also need to collaborate efficiently.
+The deep learning library often invokes collective communication library to implement collective communication primitives to transmit activations (in forward propagation) or synchronize gradients (in backward propagation) between different GPUs.
+CCL often gathers several GPU nodes as a communicator and schedules data traffic to finish a collective communication task.
+Commonly used collective communication primitives are detailed in Section II-C.
+The actual communication data traffic generated by CCL is injected into the underlying network. A variety of network protocols and topologies for distributed training are detailed in Section II-D.
 Parallelization strategy, CCL, and network are three critical components in the architecture that affect communication efficiency.
 The interplay among them and their impacts on communication efficiency are analyzed in Section II-E.
+In order to reduce communication overhead, various communication compression methods are also commonly used in practice but often sacrifice accuracy or memory.
+Communication compression methods are often associated with deep learning libraries and are out of the scope of this article due to space constraints.
 II-B Common Parallelization Strategies
-Commonly used parallelization strategies include data parallelism, model parallelism (primarily including pipeline parallelism and tensor parallelism), and emerging Mixture-of-Expert (MoE) parallelism, as illustrated by Fig.
-3, where each rectangle refers to a GPU node and each round rectangle refers to (part of) a model.
-Data parallelism is one of the most commonly used parallelization strategies whose basic idea is to distribute multiple copies of a model to different GPUs.
-Each GPU receives a subset of training data (mini-batch) during training, and gradients are averaged to update global parameters at the end of each iteration.
+Fig.
+2 shows four commonly used parallelization strategies that include data parallelism, model parallelism (includes pipeline and tensor parallelism), and emerging Mixture-of-Expert (MoE) parallelism.
+Data parallelism is one of the most commonly used parallelization strategies that distribute a subset of training data (mini-batch) to multiple copies of a model on different GPUs.
 Model parallelism, including pipeline and tensor parallelism, means splitting the model onto different GPUs.
@@
 MoE parallelism is another example of hybrid parallelism.
-The idea of data parallelism is implicit in MoE parallelism, and the expert model itself can also be split into multiple GPUs for parallel computing.
+The idea of data parallelism is implicit in MoE parallelism, and the expert model itself can also be split onto multiple GPUs for parallel computation.
 II-C Common Collective Communication Primitives
-Commonly used collective communication primitives implemented in CCL for distributed training are illustrated in Fig.
-3, where each rectangle represents a GPU node and each rounded rectangle represents a data chunk.
-Broadcast distribute data from a particular node to all other nodes, which can be used in data parallelism and some model parallelisms [8].
+Commonly used collective communication primitives implemented in CCL are illustrated in Fig. 3.
+Broadcast distributes data from a particular node to all other nodes, which can be used in data parallelism and some model parallelism [8].
 All-Gather is a many-to-many collective communication primitive where data from different nodes are distributed to all nodes.
-All-to-All transmit data among various nodes, such as data distribution in MoE parallelism [9, 10].
+All-to-All transmits data among various nodes, such as data distribution in MoE parallelism [9, 10].
 All-Reduce is a sum operation of the corresponding data chunk of each node, as shown in Fig.
@@
 Protocols and topologies are two main factors that affect network performance.
-Besides general TCP/IP protocol, distributed training often uses RDMA (such as RoCE or Infiniband) for less overhead and higher bandwidth.
+Besides general TCP/IP protocol, distributed training often uses RDMA (Remote Direct Memory Access), such as RoCE or Infiniband, for less overhead and higher bandwidth.
 The traffic flow within the network is also closely related to the network topology.
 Common network topologies used for distributed training include Fat-tree and its variants, Torus, as well as Ring and Full-Mesh topologies.
-These topologies can be combined according to practical needs.
+These topologies can be combined according to actual needs.
 For example, the NVLink topology of Nvidia’s DGX-1666https://images.nvidia.com/content/pdf/dgx1-system-architecture-whitepaper1.pdf is a combination of Ring and Full-mesh.
-II-E Communication Paradigm in the Architecture
+II-E Three-Layer Paradigm of Communication Optimization
 Parallelization strategy, CCL, and network form a three-layer paradigm of communication optimization for distributed training, as shown by the colored rectangles in Fig.
-3.
+1.
 All of them directly or indirectly affect the communication performance of the training process.
-Different parallelization strategies determine the primarily used collective communication primitives in task graphs, which affects the traffic pattern.
-Various implementations of CCL directly affect the actual network traffic.
+Different parallelization strategies determine the primarily used collective communication primitives in task graphs, which affect the traffic pattern.
+Various implementations of a communication primitive in CCL directly affect the network traffic.
 The same traffic demand often exhibits different performance under different network infrastructures.
@@
 For example, the execution process of the Ring-based All-Reduce algorithm results in a ring communication mode.
-That is, each node transmits data to its logical neighbors while all the nodes in the communicator (a communicator is a set of nodes used to implement a collective communication task) form a ring.
+That is, each node transmits data to its logical neighbors while all the nodes in the communicator (a communicator is a set of nodes used to perform a collective communication task) form a ring.
 Nvidia’s DGX-16 and Google’s Torus [4] topology both contain numerous ring structures, which are well-suited to satisfy the communication needs of Ring-based All-Reduce.
-However, layers are relatively independent in the current three-layer communication paradigm.
-Although some methods introduced in Section III involve cross-layer design, collaborative design has not yet become mainstream in general, which limits the end-to-end training performance [6].
-Therefore, we advocate a communication-efficient five-layer paradigm with additional (logical) schedulers (dashed boxes in Fig.
-3) over the current architecture.
-The inter-layer schedulers imply collaborative design, which is elaborated in Section IV.
+In Section III, we will introduce the recent advances of communication optimization in each of the three layers.
+We find that inter-layer collaboration is still limited and is a direction worth exploring.
+Therefore, we advocate a new five-layer paradigm in Section IV, which includes inter-layer schedulers to achieve more efficient communication.
 III Overview of Recent Advances
-Communication optimization in distributed training encompasses numerous aspects.
-We delve into three of the most pertinent elements: parallelization strategy, CCL, and network, which make up the current three-layer communication paradigm.
-We use exposed communication to denote the communication time that cannot overlap with computation time.
-In this section, our focus is confined to the contributions of a specific study pertaining to communication optimization, that is, how to reduce exposed communication in distributed training.
+| Specific works | Layer | Focus | How to reduce exposed communication | Time for decision | Scala- bility | Beneficial effect | Deployability |
+| Megatron-lm [7] | Para. | Intra-layer model parallelism | Reduce traffic by removing a synchronization point | Offline | High | 74% linear scaling on 512 GPUs | Integrated in PyTorch |
+| Alpa-Comm [8] | “Cross-mesh resharding” problem | Reduce inter-node traffic and overlap with computation | Offline & Online | Mid | 1.1x training speedup on GPT | 900 lines of C++ and 2.5K lines of python | |
+| PTD-P [1] | Combination of different parallelism | Overlap with computation by a interleaved pipeline | Offline | High | 502 petaFLOP/s on 3072 GPUs | Integrated in PyTorch | |
+| Lina [9] | All-to-All bottleneck in MoE parallelism | Overlap All-Reduce comm. with computation | Offline | Mid | Up to 1.73x training speedup | Integrated in PyTorch | |
+| Janus [10] | Data-centric paradigm for MoE parallelism | Reduce All-to-All traffic and overlap with computation | Offline & Online | High | Up to 2.4x training speedup | Integrated in PyTorch | |
+| NCCL | CCL | Comm. primitive implementation | Select efficient communication primitive algorithms | Offline & Online | High | Widely used infrastructure | Integrated in PyTorch and TensorFlow |
+| Blink [11] | Comm.
+primitive optimization for underlying topology | Overlap flows by fine-grained scheduling to improve link utilization and shorten primitive completion time | Online | Low | Up to 1.4x training speedup | Integrated in PyTorch and TensorFlow | |
+| SCCL [12] | Offline | Low | 1.14x to 2.2x training speedup | Integrated in MSCCL (Microsoft CCL) | |||
+| TACCL [5] | Synthesizing comm. primitive with human sketch | Offline | Mid | Up to 2.36x training speedup on BERT | Integrated in PyTorch | ||
+| SYNDI-CATE [13] | Co-optimization of runtime schedule and execution | Overlap primitives to promote link utilization | Offline & Online | Mid | 1.21x to 1.74x training speedup | Integrated in PyTorch | |
+| TPUv4 [4] | Net. | Supercomputer for machine learning | Efficient topology for distributed machine learning | Offline | High | Nearly 10x faster overall | Up to 4096 chips connected by OCSs |
+| TopoOpt [2] | Co-optimization for topology and parallelization strategy | Match topology with parallelization strategy | Offline | High | Up to 3.4x training speedup | With FlexFlow’s training engine and modified NCCL |
+Communication optimization for distributed training encompasses numerous aspects.
+We delve into three of the most pertinent elements: parallelization strategy, CCL, and network.
+In this section, our focus is confined to the contributions of a specific study pertaining to communication optimization, that is, how to reduce exposed communication (the communication time that cannot overlap with computation time) in distributed training.
 Table I provides an overview of the research advances with some primary concerns.
-Layer |
-Main focus |
-How to reduce exposed communication |
-Cross-layer co-design |
-Time for decision |
-Implementation |
-Method |
-Beneficial effect |
-Megatron-lm [7] |
-Model Parallelism |
-Efficient intra-layer model parallel strategy |
-An efficient tensor partitioning strategy |
-Reducing the amount of traffic by removing a synchronization point |
-Offline |
-Insert a few communication operations in naive pytorch |
-AlpaComm [8] |
-Addressing “cross-mesh resharding” problem in model parallelism |
-Broadcast based resharding method and eager 1F1B pipeline schedule |
-Reducing inter-node traffic and overlap communication with computation |
-With CCL |
-On/Offline |
-900 lines of C++ and 2.5K lines of python |
-PTD-P [1] |
-Hybrid Parallelism |
-Combination of pipeline, tensor and data parallelisms |
-A interleaved pipelining schedule |
-Overlapping communication and computation in pipeline parallelism |
-Offline |
-An extension to Megatron-LM codebase |
-Lina [9] |
-MoE Parallelism |
-All-to-All bottleneck in MoE parallelism |
-Priority scheduling All-to-All with All-Reduce micro-ops |
-Overlapping All-Reduce communication with computation |
-With CCL |
-Offline |
-Based on DeepSpeed MoE with 7500 lines of code |
-Janus [10] |
-A data-centric paradigm for MoE parallelism |
-Move experts to reduce All-to-All traffic when benificial and a provident prefetch strategy |
-Reducing All-to-All traffic and overlap communication with computation |
-On/Offline |
-A plugin of 4.3K lines of code in PyTorch |
-NCCL |
-Collective Communi- cation Library (CCL) |
-Multi-GPU and multi-node communication primitives optimized for NVIDIA hardware |
-Select communication primitive algorithms such as DBTree and Ring for All-Reduce, etc. |
-Mainly accelerating All-Reduce communication |
-For particular hardware |
-On/Offline |
-Integrated in Deep learning frameworks such as PyTorch and TensorFlow. |
-Blink [11] |
-Synthesize collective communication algorithms for underlying topology |
-Packing spanning trees with MWU and ILP algorithms |
-Overlap flows within primitives to promote link utilization |
-For topology |
-Online |
-Seamlessly plugged into pytorch and tensorflow |
-SCCL [12] |
-Synthesize collective communication algorithms for underlying topology |
-Synthesizing pareto-optimal algorithms with SMT encoding |
-Overlap flows within primitives to promote link utilization |
-For topology |
-Offline |
-Evaluated on different 8-GPU systems such as Nvidia DXG-1 and Gigabyte Z52 |
-TACCL [5] |
-Synthesize collective communication algorithms for topology and training task |
-Routing (MILP), ordering (Greedy algorithm) and scheduling (MILP) with communication sketches |
-Overlap flows within primitives to promote link utilization |
-For topology |
-Offline |
-Extension of NCCL with TACCL runtime as interpreter |
-SYNDI-CATE [13] |
-Co-optimize collective communication algorithms with runtime schedule |
-MCMC Search with termination |
-Overlap primitives to promote link utilization |
-With execution plan |
-On/Offline |
-UCC library interface (pytorch integrated) |
-TPUv4 [4] |
-Network |
-Google’s third supercomputer specific for ML |
-Architectural |
-Up to 4096 chips connected by OCSs |
-TopoOpt [2] |
-Co-optimizing computation, communication and network topology |
-A group theory-inspired algorithm called TotientPerms |
-The best network topology and routing plan, together with a parallelization strategy |
-With parallelization strategy |
-Offline |
-A 12-node prototype with FlexFlow’s training engine and modification to NCCL |
+In the table, we use Para., CCL and Net.
+to represent Parallelization Strategy, Collective Communication Library and Network respectively.
+We evaluate the studies from many aspects such as time for decision, scalability, and deployability.
+Time for decision indicates whether it plays a role during training (online) or before training (offline), which is crucial for practical applications.
+Scalability refers to the ability to be applied to large scenarios.
+We also show representative performance metrics under specific experimental settings in beneficial effect.
 III-A Parallelization Strategy
-The communication overhead in distributed training mainly arises from parallelization.
-Different implementations of parallelization strategies affect the communication pattern and network traffic during the model training process [2].
+The communication overhead of distributed training stems from the data synchronization requirements brought about by parallelization.
+Different parallelization strategies determine different communication types, which in turn bring about different network traffic patterns [2].
 The most straightforward and most widely used parallelization strategy is data parallelism.
 As mentioned in Section II, multiple copies of the same model are replicated across different nodes in data parallelism.
-Each iteration requires all nodes to synchronize gradients, which are used for global parameter updates.
+Each iteration requires all nodes to synchronize gradients, which are used to update global parameters.
 Consequently, data parallelism results in a significant amount of All-Reduce communication.
-To train a large-scale model that can’t fit into the memory of a single GPU, researchers introduced model parallelism, which includes pipeline and tensor parallelism.
-Tensor parallelism divides models within a layer and requires intensive communication in training.
-Researchers from Nvidia proposed an efficient “intra-layer parallelism” approach called Megatron-lm [7] to reduce the communication traffic in tensor parallelism.
-Pipeline parallelism necessitates only the transfer of activations/gradients between adjacent nodes in the logical pipeline, which is a point-to-point communication pattern that significantly reduces the communication cost compared to data parallelism and tensor parallelism.
-Because overlapping communication and computation is hard to achieve in naive pipeline parallelism, an interleaved pipeline scheduling strategy [1] is proposed where each device is assigned multiple data chunks to gain the chance of communication-computation overlap.
+However, due to memory constraints, very large models cannot be trained using only data parallelism.
+To address this, researchers introduced model parallelism, which includes pipeline and tensor parallelism.
+Tensor parallelism divides a layer of a model, requiring intensive communication for synchronizing activations/gradients at each layer during propagation.
+To reduce the communication traffic in tensor parallelism, researchers from Nvidia proposed an efficient “intra-layer parallelism” approach called Megatron-lm [7].
+It achieved 77% of linear scaling with a 8.3 billion parameters model on 8 GPUs, and 74% of linear scaling on 512 GPUs.
+To further divide the model by layers, researchers introduced pipeline parallelism.
+Pipeline parallelism only requires transferring activations/gradients between adjacent nodes in the logical pipeline, which is a point-to-point communication mode that requires less communication compared to data parallelism and tensor parallelism.
+Since communication and computation are difficult to overlap in the original pipeline, researchers introduced an interleaved pipeline scheduling strategy [1] where each device is assigned multiple data chunks at a time to gain the chance of communication-computation overlap.
 However, pipeline parallelism, as an “inter-layer parallelism” approach, is not without its flaws.
@@
 Various parallelization strategies are often used in conjunction.
-For example, PTD-P [1] not only combines Megatron-lm’s Tensor parallelism and pipeline parallelism but also introduces data parallelism to scale to thousands of GPUs.
-AlpaComm [8] identifies the “cross-mesh resharding” problem that arises when combining tensor and pipeline parallelism and proposes a communication optimization method based on broadcasting and an “overlapping-friendly” pipeline scheduling scheme to accelerate end-to-end training.
-As an emerging model structure, the MoE model is gradually gaining popularity due to its fast training/inference speed.
-The training process of the MoE model selects several appropriate experts for a data sample, which inherently has good compatibility with parallelization.
-MoE parallelism can be achieved by simply dispersing the “experts” on different GPUs, introducing All-to-All traffic when data is distributed between different GPUs.
-Additionally, the model parameters of the non-expert part are globally shared in MoE parallelism, so there will also be All-Reduce traffic.
-Lina [9] uses a scheduling strategy that prioritizes All-to-All traffic and further splits the All-Reduce tasks to overlap with computation as much as possible, thus reducing exposed communication and achieving training acceleration.
+For example, PTD-P [1] combines tensor parallelism, pipeline parallelism and data parallelism to scale to thousands of GPUs.
+It can train a model with 1 trillion parameters at 502 petaFLOP/s on 3072 GPUs (per-GPU throughput of 52% of theoretical peak).
+AlpaComm [8] identifies the “cross-mesh resharding” problem that arises when combining tensor and pipeline parallelism.
+It proposes a communication optimization method based on broadcasting and an “overlapping-friendly” pipeline scheduling scheme to accelerate end-to-end training.
+It shows a 1.1x speedup compared to their state-of-the-art baseline because of overlapping.
+The MoE model is an emerging model structure that is gaining popularity due to its high performance, flexibility and fast training/inference speed.
+During the training process, each data sample is assigned to several appropriate experts, which inherently has good compatibility with parallelization.
+MoE parallelism can be achieved by simply dispersing the “experts” on different GPUs.
+This introduces All-to-All traffic when data is distributed between different GPUs.
+Additionally, non-expert part of the model will also introduce All-Reduce traffic during training.
+There are many studies on optimizing MoE parallelism.
+Lina [9] uses a scheduling strategy that prioritizes All-to-All traffic and further splits the All-Reduce tasks to overlap with computation as much as possible, thus reducing exposed communication.
 Janus [10] studies MoE parallelism acceleration from a new perspective.
 Since the size of expert models and training datasets varies, Janus innovatively proposes a “data-centric” model that “moves experts instead of data”.
-This significantly reduces the amount of communication during the training process under certain conditions (e.g., when the scale of experts is smaller than the scale of data), thus accelerating the training process.
+This significantly reduces the amount of communication traffic up to 16x under certain conditions (e.g., when the scale of experts is smaller than the scale of data) and achieves up to 2.06× speedup compared with MoE training system at the time.
 III-B Collective Communication Library
 Communication tasks of distributed training are often implemented by CCL.
-The current predominant NVIDIA Collective Communication Library (NCCL)777https://developer.nvidia.com/nccl dynamically selects established primitive algorithms based on different situations.
-For instance, NCCL selects the Ring and Double Binary Tree algorithms for All-Reduce 888https://developer.nvidia.com/blog/massively-scale-deep-learning-training-nccl-2-4 for different workloads and topologies to accelerate the training process.
+The current predominant NVIDIA Collective Communication Library (NCCL)777https://developer.nvidia.com/nccl dynamically selects established algorithms based on different situations.
 However, NCCL’s optimization for specific hardware underscores the need for a more universal and adaptable CCL to speed up distributed training across different hardware and topologies.
-Therefore, many researchers took different approaches to implement communication primitives more efficiently.
+Therefore, many researchers take different approaches to implement communication primitives more efficiently.
 For example, Blink [11] dynamically generates optimal communication primitives by packing spanning trees rather than writing and optimizing them manually.
-Blink uses integer linear programming (ILP) to reduce the number of trees and improve performance, and also leverages heterogeneous communication channels to utilize full bandwidth.
-At the time, Blink achieved up to 40% speedup compared to NCCL in end-to-end DNN training iterations.
-SCCL [12] takes a more unified approach to achieve efficiency: automatically synthesizing high-performance communication primitives for a given topology.
-To achieve this, SCCL designs a cost model to evaluate the latency and bandwidth cost of algorithms and then searches for algorithms on the Pareto frontier.
-However, SCCL also faces problems of high complexity.
-The communication primitive algorithm generation problem is encoded into a Mixed Integer Linear Program (MILP), which is unfortunately NP-hard.
-To reduce the complexity of generating collective communication algorithms, TACCL [5] is a representative work that introduces a “human-in-the-loop” approach, whose paradigm is shown in Fig.
+It uses integer linear programming (ILP) to reduce the number of trees and improve performance, and also leverages heterogeneous communication channels to fully utilize bandwidth.
+At the time, Blink achieved up to 1.4x speedup compared to NCCL in end-to-end DNN training iterations.
+SCCL [12] takes a more unified approach to automatically synthesize high-performance communication primitives for a given topology.
+It designs a cost model to evaluate the latency and bandwidth cost of algorithms and then searches for algorithms on the Pareto frontier.
+SCCL is 1.14x to 2.2x faster than NCCL’s All-Gather on different sizes of data at the time.
+However, generating communication primitives is a Mixed Integer Linear Program (MILP) problem, which is unfortunately NP-hard and can require a long time to solve.
+Although Blink and SCCL focus on optimization within small heterogeneous topologies and might not be practical to scale to large-cluster scenarios with tens of thousands of GPUs because of complexity, they lay the groundwork for future CCL optimization works.
+Based on SCCL, TACCL [5] introduces a “human-in-the-loop” approach to reduce the complexity of generating collective communication algorithms.
+The workflow of TACCL is shown in Fig.
 4.
@@
 Compared with NCCL, TACCL achieves up to 2.36x end-to-end training speedup for BERT and 1.94x for Transformer-XL.
-Different from previous methods focused on speeding up single primitives, SYNDICATE [13] proposes a novel abstraction to overlap primitives in a fine-grained pattern, breaking large communication operations into smaller units, thus allowing higher flexibility.
-SYNDICATE then develops a joint optimizer based on Markov Chain Monte Carlo (MCMC) search to optimize the joint action of the control plane and the data plane.
+Different from previous methods focused on speeding up single primitives, SYNDICATE [13] proposes a novel abstraction to overlap primitives in a fine-grained pattern.
+It breaks large communication operations into smaller units, thus allowing higher scheduling flexibility.
+SYNDICATE jointly optimizes the implementation and scheduling of communication primitives, enabling parallel execution of primitives.
+It improves the speed of training state-of-the-art large models by 21-45%.
+SYNDICATE embodies the idea of cross-layer collaborative design, but is still a preliminary exploration.
 III-C Network
 Distributed training focuses on GPU-to-GPU communication, which needs to consider intra/inter-host communication protocols.
-Common intra-host communication protocols include the general PCIe protocol and Nvidia’s NVLink/NVSwitch.
-Inter-host communication protocols include traditional TCP/IP and more efficient RDMA (e.g., RoCE and Infiniband).
-Based on certain communication protocols, the underlying network topology also significantly influences the communication efficiency of distributed training.
-For instance, as mentioned in Section II, Nvidia utilizes a Ring and fully connected Mesh topology in their DXG series, and Google employs a 3D-Torus topology in their TPUv4 [4].
+Common intra-host communication protocols include the general PCIe protocol and Nvidia’s NVLink/NVSwitch, while inter-host protocols include traditional TCP/IP and more efficient RDMA (e.g., RoCE and Infiniband).
+Besides protocols, the underlying network topology also significantly influences the communication efficiency of distributed training.
+For instance, as mentioned in Section II, Nvidia utilizes a Ring and fully-connected Mesh topology in their DGX series, and Google employs a 3D-Torus topology with optical circuit switches (OCSs) in their TPUv4 [4].
 These topologies contain a large number of ring structures, which enhance the efficiency of Ring-based All-Reduce and Point-to-Point communications (such as pipeline parallelism or broadcast-based hybrid parallelism [8]).
-However, static topologies can rarely adapt to dynamically changing job requirements, leading researchers to propose TopoOpt [2].
+However, static topologies can rarely adapt to dynamically changing job requirements, motivating researchers to propose TopoOpt [2].
 TopoOpt takes advantage of the reconfigurability of optical switches, enabling a scheme that co-optimizes the topology and parallelization strategy.
 Nevertheless, current optical switching technology still has issues of high reconfiguration latency, making it challenging to perform dynamic topology adjustments between iterations.
-As such, TopoOpt remains a relatively coarse-grained optimization, that is, reconfiguring the interconnection between training servers of each job only before the job starts and keeps the topology until the training is complete.
-Anyway, research on communication optimization for distributed training from a network perspective is still insufficient, which is prospected in Section IV.
-III-D Collaboration Design in Current Advances
-As observed in Table I, the research advances mentioned above are primarily limited to optimizations within a layer, with less cross-layer co-design.
+As such, TopoOpt remains a relatively coarse-grained optimization.
+It reconfigures the topology only before the job starts and maintains it until the training is complete.
+In conclusion, research on communication optimization for distributed training from a network perspective is still insufficient, which requires further exploration.
+III-D Collaborative Design in Current Advances
 Current advances involving cross-layer optimization can be categorized into two types.
-The first type is passive collaborative design.
+The first is passive collaborative design.
 For instance, AlpaComm [8] and Lina [9] optimize communication for a specific parallelization strategy, necessitating modifications to communication primitives.
-Another example is that generative CCL (such as Blink [11], SCCL [12] and TACCL [5]) has to be aware of network topologies because their main significances are to customize collective communication algorithms for specific topologies.
-The second type of work involves proactive attempts at collaborative optimization yet remains exploratory.
+Another example is that generative CCL (such as Blink [11], SCCL [12] and TACCL [5]) has to be aware of network topologies because their main focuses are to customize collective communication algorithms for specific topologies.
+The second type is proactive collaborative design.
 For example, SYNDICATE [13] jointly optimizes the implementation and scheduling of communication primitives, allowing parallel execution of primitives.
 TopoOpt [2] collaboratively optimizes network topology and parallelization strategies, obtaining optimized combination strategies through alternating optimization techniques.
+However, these works are still preliminary attempts and require further research.
 In summary, collaborative communication optimization in distributed training is still a domain that needs further exploration.
-Section IV details the importance of co-design and potential research opportunities.
+Section IV details the importance of collaborative design and potential research opportunities.
 IV Opportunities
-As noted in Section III, collaborative design has not received widespread attention in current research advances.
-However, different from general tasks in data centers, the communication traffic patterns in distributed training are known in most cases [6, 14], providing motivations and opportunities for cross-layer collaboration.
-Furthermore, the optimization objectives for communication tasks are no longer the general-purpose flow complete time (FCT) but rather the job complete time (JCT).
-In this context, communication tasks are merely a component of the entire distributed training process.
-Optimizing JCT necessitates considering the dependency between communication and computation instead of just optimizing each communication task.
-Communication tasks should be optimized to minimize computation blocking time, which can only be achieved through “cross-layer” scheduling to meet this new optimization objective.
-To this end, we advocate a communication-efficient five-layer paradigm with (logical) schedulers as the middleware, as illustrated in Fig.
-5.
-The schedulers not only allow lower-layer facilities to perceive the requirements of higher-layer applications but also enable higher-layer applications to be aware of the capabilities of lower-layer facilities.
-We refer to this as “Vertical” co-design.
-The schedulers can also perform global scheduling across multiple jobs, achieving “Horizontal” collaborative optimization.
-Considering the heterogeneous topologies of distributed training and the development of emerging technologies such as in-network aggregation, we also provide a prospect for “Intra-Inter” and “Host-Net” co-design.
-“Vertical” co-design: Refers to exchanging necessary information between layers to optimize communication efficiency.
-For example, CCL can acquire global traffic demand from the parallelization strategy and generate the optimal communication primitive algorithms in conjunction with the network state.
+Different from general communication tasks in data centers, the data traffic patterns of distributed training are predictable in most cases [6, 14].
+Furthermore, the optimization objectives of communication tasks are no longer the general-purpose flow completion time (FCT) but rather the job completion time (JCT).
+Therefore, optimizing JCT necessitates considering the dependency between communication and computation tasks and the relationship between different communication tasks instead of just optimizing a single communication task.
+These provide motivations and opportunities for collaborative design to optimize efficiency.
+In this section, we use a simple case study (Fig.
+5(b)) to illustrate the necessity of collaborative design in distributed training.
+We provide research opportunities from two aspects: the communication-efficient five-layer paradigm and the collaboration of heterogeneous resources.
+IV-A Communication-Efficient Five-Layer Paradigm
+In Section II, we have introduced the current three-layer paradigm of communication optimization for distributed training, which includes Parallelization strategy, CCL, and Network.
+In order to promote cross-layer collaboration, we extend the three-layer communication paradigm to a communication-efficient five-layer paradigm, as illustrated in Fig.
+5(a), where the red arrows indicate possible information exchange for collaborative optimization.
+In the new paradigm, the Parallelization Strategy (Para.), CCL, and Network (Net.) are still the three main layers, while the (logical) schedulers are added as the middleware.
+The task scheduler is to schedule the communication tasks required by the parallelization strategy, and the flow scheduler is to schedule the communication flows generated by CCL.
+Standardized interfaces and protocols are required to facilitate seamless communication between layers, which allow lower-layer facilities to perceive the requirements of higher-layer applications and enable higher-layer applications to be aware of the capabilities of lower-layer facilities.
+Each layer can use shared information to optimize with a broader perspective, achieving “Vertical” co-design.
+For example, CCL can generate task-specific communication primitive algorithms based on the global traffic demand from the parallelization strategy and the network state, thus minimizing JCT.
 Furthermore, cross-layer schedulers can be constructed to strategically schedule communication tasks based on traffic load or optimize task graphs according to the traffic patterns of relevant communication primitives.
-Echelon [14] is a flow scheduling work that embodies “Vertical” co-design, where novel communication optimization objectives are designed based on the dependency between communication and computation under various parallelization strategies, aiming to optimize communication towards minimizing JCT.
-Another example of “Vertical” co-design is that network can also use traffic demand for optimization in distributed training.
-As the general infrastructure, the network is often unaware of traffic demand.
-This is natural in common data center scenarios, where the network can only implement passive adjustments such as RTT-based congestion control due to the random nature of flow arrivals.
-However, traffic patterns are often predictable in distributed training scenarios[6].
-Consequently, the network itself can also proactively adjust by taking traffic demands of high-layer training tasks into account, thus enhancing goodput.
-This in turn reduces communication time and accelerates the distributed training process.
-“Horizontal” co-design: Primarily focuses on the coordination among different training jobs or various communication tasks.
-Within a single training job, the traffic from multiple concurrent collective communication primitives competes for network resources.
-Similarly, in the multi-job scenario, the traffic from numerous concurrent training jobs also competes for network resources.
-Consequently, optimization strategies tailored to a single communication primitive or training job may suffer varying degrees of inefficacy.
+In the case study of Fig.
+5(b), two concurrent flows of Job1 compete for ToR switch at ①, which may lead to congestion.
+However, the “deadline” of the two flows may be different, and the scheduler can adjust the priority of the flows to avoid blocking the subsequent computation process.
+In other words, the scheduler can consider both the task graphs and the network constraints to optimize the communication tasks, which is an example of “Vertical” co-design.
+This collaborative schedule can minimize JCT for the entire Job1, not only FCT for a certain flow.
+Echelon [14] is a flow scheduling example that embodies “Vertical” co-design, where novel communication optimization objectives are designed based on the dependency between communication and computation under various parallelization strategies, aiming to optimize communication towards minimizing JCT.
+The schedulers can also perform global scheduling across multiple jobs, achieving “Horizontal” co-design.
+In multi-job scenarios, the traffic from numerous concurrent training jobs competes for network resources.
+As shown in Fig.
+5(b), flows of Job1 and Job2 compete for the same ToR switch at ②.
+Consequently, optimization strategies tailored to a single training job may suffer varying degrees of inefficacy.
 Therefore, it is imperative to employ horizontal co-design at different granularities to fully utilize network resources, ultimately reducing JCT of training tasks.
+A potential solution is to implement dynamic resource manager that allocates resources based on real-time demands.
 CASSINI [6] attempts horizontal co-design through the perspective of multi-task scheduling.
-It analyses the consequence of non-coordinated job scheduling in multi-job scenarios, highlighting the potential traffic conflicts arising from the periodic patterns of network resource utilization by different training jobs.
-Subsequently, CASSINI introduces a design for “staggering peak” scheduling for different jobs, achieving high utilization of network resources, thus enhancing the performance of multiple training jobs.
-“Intra-Inter” co-design: Refers to the coordinated utilization of multidimensional network resources, primarily including the “intra-host” and “inter-host” network.
-The current networks used in distributed training exhibit substantial heterogeneity.
-The interconnection patterns between different GPUs within the same host include PCIe or NVLink, while those between different hosts include RoCE or Infiniband, etc.
-Given that GPU (rather than host) is the minimum unit in the communication tasks of distributed training, two modes of GPU-to-GPU communication exist: within the same host and across different hosts.
+It performs “staggering peak” scheduling for different jobs, which improves the utilization of network resources.
+In addition, horizontal collaboration also enables multi-job Service Level Agreement (SLA) management.
+IV-B Collaborate Design for heterogeneous resources
+Considering the heterogeneous network topologies of distributed training and the development of emerging technologies such as in-network aggregation, we provide a prospect for “Intra-Inter” and “Host-Net” co-design.
+“Intra-Inter” co-design refers to the coordinated utilization of multidimensional network resources, primarily including the “intra-host” and “inter-host” network.
+GPUs are typically interconnected via fast PCIe or NVLink within a host, while inter-host connections are established through relatively slower RoCE or InfiniBand.
 In this context, the issue of heterogeneous link bandwidths within and between hosts is inevitably confronted.
-An example of “Intra-Inter” co-design is that TACCL [5] utilizes a profiling approach to perceive the characteristics of heterogeneous links, which serves as a basis for optimizing communication primitive algorithms.
-“Host-Net” co-design: Primarily refers to the collaboration of computing resources between the end-host and programmable switches within a network.
+Let us suppose that Flow1 of Job1 is to implement a Ring All-Reduce among GPUs across two hosts in Fig.
+5(b).
+Then the bottleneck of this communication task will be the bandwidth of the inter-host link, which is slower than the intra-host link.
+Solving this problem requires real-time sensing and scheduling of heterogeneous link resources.
+TACCL [5] utilizes a profiling approach to perceive the characteristics of heterogeneous links in order to optimize the communication primitive algorithms, which embodies the concept of “Intra-Inter” co-design.
+“Host-Net” co-design primarily refers to the collaboration of computing resources between the end-host and programmable switches within a network.
 In recent years, with the development of programmable switch technology, in-network aggregation has gradually become a new computing paradigm.
-In-network aggregation not only fully utilizes the computing capabilities of programmable switches within the network but also reduces the traffic running in the network, achieving multiple purposes in one stroke.
+In-network aggregation uses the computing capabilities of programmable switches within the network to reduce the traffic running in the network.
+As shown by the blue line in Fig.
+5(b), the two flows of Job2 are aggregated into one flow at the ToR switch, thus reducing further communication overhead.
 However, since existing protocols are designed for end-to-end communication, implementing in-network aggregation usually requires the support of new protocols.
@@
 This approach also reduces the network traffic for a collective communication task.
-ATP embodies the concept of “Host-Net” co-design and represents an exemplary work of in-network computing in the field of distributed training.
+ATP embodies the concept of “Host-Net” co-design.
 Nevertheless, in-network aggregation has not been widely applied in the production environment due to its complexity.
 “Host-Net” co-design remains a direction worthy of long-term research.
+IV-C Open Issues
+Cross-layer information exchange mechanism: The lack of effective information exchange mechanism across different layers makes collaborative optimization difficult.
+How to design good protocols to achieve efficient information exchange between different layers is a problem worth studying.
+Utilization of heterogeneous network resources: The full utilization of heterogeneous network resources is a long-term complex problem, involving the awareness of the real-time status of different network resources and timely scheduling.
+This poses high requirements on the complexity and accuracy of the scheduler.
+Dynamic optimization of communication libraries: Current communication optimization strategies are usually offline, and existing generative CCLs are difficult to meet the real-time requirements of online operation.
+However, network status is dynamic, and communication patterns also change in emerging model structures such as MoE.
+Therefore, the dynamic optimization of communication libraries is an important research direction.
+Network topology specific for distributed training: Current network topologies are usually optimized for general tasks (e.g., fat-tree) or common communication primitive implementations (e.g., Ring-based All-Reduce).
+How to design or generate an efficient network topology for distributed training tasks and introduce dynamic adjustments to the topology according to the change in traffic load is a research direction worth exploring.
+Communication Optimization with emerging techniques: With the emergence of new hardware (such as programmable switches, optical switches, etc.), it is important to study the impact of these new hardware on communication optimization and explore how to make full use of these hardware to improve communication efficiency.
 V Conclusion
 Due to the dramatic increase in the number of parameters of large-scale deep neural network models, it is imperative to develop effective communication optimization methods in distributed training.
-In this article, we first provide a general distributed training architecture and the current three-layer paradigm of communication optimization consisting of parallelization strategy, CCL, and network.
-We then review the latest advances in the current three-layer paradigm.
-Furthermore, we analyze the shortcomings of the current paradigm and advocate a communication-efficient five-layer paradigm, highlighting great opportunities for “cross-layer” co-design.
-We hope to shed some light on future research on communication optimization for distributed training.
+In this article, we first introduce a general distributed training architecture from the communication perspective and summarize the current three-layer paradigm of communication optimization, including parallelization strategy, CCL, and network.
+We review the latest advances in the current three-layer paradigm and summarize their important features in a table.
+Then, we look forward to future research directions from the perspective of collaborative design.
+We advocate a communication-efficient five-layer paradigm and collaborative design for heterogeneous resources.
+Finally, we propose some open issues.
+We believe that this article can shed some light on future research on communication optimization for distributed training.
+Acknowledgments
+This work was supported by NSFC Project under Grant 62132009 and Grant 62221003.